# Visual Explanation via Similar Feature Activation for Metric Learning

### v2

@@ 3.1 Channel-wise Contribution Importance Score
Let $f(\cdot)$ to denote a CNN backbone network for metric learning. $Q$ and $S^{c}$ denote a query image and a support image from the specific class $c$, respectively. Thus, $f(Q)$ and $f(S^{c})$ represents the feature map from the query image and the support image respectively, so $f(Q)$, $f(S^{c})$ $\in R^{N\times H\times W}$ where NN, $H$, and $W$ denote the number of channels, height and width of the feature map respectively. Let $p(\cdot)$ to denote a pooling function. The query representation $V^{Q}$ (feature vector) can be denoted by $p(f(Q))\in R^{N}$. Suppose that there are $K$ support images from the class $c$, the support representation $V^{S}$ can be expressed by
@@
$V^{Q}$ and $V^{S}$ represent the feature vectors of NN elements from the query image $Q$ and the support image $S$ respectively. Let $V^{Q}=[v^{Q}_{1},v^{Q}_{2},\cdots,v^{Q}_{N}]$ and $V^{S}=[v^{S}_{1},v^{S}_{2},\cdots,v^{S}_{N}]$, the scalar $v^{Q}_{n}$ and $v^{S}_{n}$ are the elements at the $n$-th channel of $V^{Q}$ and $V^{S}$ respectively and $n=1,\cdots,N$. The $n$-th channel CIS $\omega_{n}$} can be calculated by the following equation,
@@
where NN denotes the number of channels of the feature map out of the final layer from a CNN backbone.
@@ 3.3 Visual Explanation for CNN using Cosine Similarity
$+$ $\frac{\hat{v}^{Q}_{n}\cdot\hat{v}^{S}_{n}}{(N-1)(2-2\text{cos}(V^{Q},V^{S}))}.$
@@
Fig. 3: The qualitative comparison between the proposed SFAM and 33 baseline methods for FRN (ResNet12) that uses Euclidean distance as the similarity metric in 55-way 11-shot image classification task.
@@ 4.2 Qualitative Experiment
To evaluate the effectiveness of the proposed SFAM, the qualitative comparison between SFAM and 44 explanation methods (e.g.,Decomposition, RAM, CG-RAM, and Grad-FAM) are conducted on the same images for FRN (ResNet12) and GPW (ResNet50). It is worth noting that Decomposition fails to interpret CNN model that using Euclidean Distance as the metric. Fig. 3 illustrates the comparison of 44 images for FRN (ResNet12) in 55-way 11-shot classification. As shown in Fig. 3, the proposed SFAM can correctly highlight the similar features (the birds’ body between the left query and left support images and the bird’s head between the right query and support images). However, the regions highlighted by RAM, GradFAM, and CG-RAM are not the similar features between two birds actually. Fig. 4 visualizes the comparison of 44 samples for GPW (ResNet50) in image retrieval. As shown in Fig. 4, the proposed SFAM can successfully generate the explanation maps for metric-learning CNN model using cosine as the similarity metric.
@@ 4.3 Quantitative Experiment
To verity the localization capacity of the proposed SFAM, the quantitative experiments are conducted on FRN (ResNet12) and GPW(ResNet50) to compare our method with 44 baseline methods (e.g., Decomposition, RAM, CGRAM, and GradFAM). IoU(intersection over Union) [GradCAM, LayerCAM, Decomposition] and localization accuracy (IoU $\geq 0.5$[LayerCAM, Decomposition] are adopted to measure the localization capacity. The threshold is set to $0.2$ in our experiment. Average IoU is calculated over all the correctly predicted samples. Specifically, in few-shot image classification, 55-way 11-shot is chosen for comparison, the average IoU is calculated over $10,000$ episodes. Table 1 and Table 2 show the quantitative comparison result. From the results, the proposed SFAM achieves the best performance in average IoU and localization accuracy, outperforming other methods by at least $2.2\%$.

### v4

@@ 3.1 Channel-wise Contribution Importance Score
Let $f(\cdot)$ to denote a CNN backbone network for metric learning. $Q$ and $S^{c}$ denote a query image and a support image from the specific class $c$, respectively. Thus, $f(Q)$ and $f(S^{c})$ represents the feature map from the query image and the support image respectively, so $f(Q)$, $f(S^{c})$ $\in R^{N\times H\times W}$ where $N$, $H$, and $W$ denote the number of channels, height and width of the feature map respectively. Let $p(\cdot)$ to denote a pooling function. The query representation $V^{Q}$ (feature vector) can be denoted by $p(f(Q))\in R^{N}$. Suppose that there are $K$ support images from the class $c$, the support representation $V^{S}$ can be expressed by
@@
$V^{Q}$ and $V^{S}$ represent the feature vectors of $N$ elements from the query image $Q$ and the support image $S$ respectively. Let $V^{Q}=[v^{Q}_{1},v^{Q}_{2},\cdots,v^{Q}_{N}]$ and $V^{S}=[v^{S}_{1},v^{S}_{2},\cdots,v^{S}_{N}]$, the scalar $v^{Q}_{n}$ and $v^{S}_{n}$ are the elements at the $n$-th channel of $V^{Q}$ and $V^{S}$ respectively and $n=1,\cdots,N$. The $n$-th channel CIS $\omega_{n}$} can be calculated by the following equation,
@@
where $N$ denotes the number of channels of the feature map out of the final layer from a CNN backbone.
@@ 3.3 Visual Explanation for CNN using Cosine Similarity
$+$ $\frac{\hat{v}^{Q}_{n}\cdot\hat{v}^{S}_{n}}{(N-1)(1-\text{cos}(V^{Q},V^{S}))}.$
@@
Fig. 3: The qualitative comparison between the proposed SFAM and 33 baseline methods for FRN (ResNet12) that uses Euclidean distance as the similarity metric in $5$-way 11-shot image classification task.
@@ 4.2 Qualitative Experiment
To evaluate the effectiveness of the proposed SFAM, the qualitative comparison between SFAM and $4$ explanation methods (e.g.,Decomposition, RAM, CG-RAM, and Grad-FAM) are conducted on the same images for FRN (ResNet12) and GPW (ResNet50). It is worth noting that Decomposition fails to interpret CNN model that using Euclidean Distance as the metric. Fig. 3 illustrates the comparison of $4$ images for FRN (ResNet12) in $5$-way 11-shot classification. As shown in Fig. 3, the proposed SFAM can correctly highlight the similar features (the birds’ body between the left query and left support images and the bird’s head between the right query and support images). However, the regions highlighted by RAM, GradFAM, and CG-RAM are not the similar features between two birds actually. Fig. 4 visualizes the comparison of $4$ samples for GPW (ResNet50) in image retrieval. As shown in Fig. 4, the proposed SFAM can successfully generate the explanation maps for metric-learning CNN model using cosine as the similarity metric.
@@ 4.3 Quantitative Experiment
To verity the localization capacity of the proposed SFAM, the quantitative experiments are conducted on FRN (ResNet12) and GPW(ResNet50) to compare our method with $4$ baseline methods (e.g., Decomposition, RAM, CGRAM, and GradFAM). IoU(intersection over Union) [GradCAM, LayerCAM, Decomposition] and localization accuracy (IoU $\geq 0.5$[LayerCAM, Decomposition] are adopted to measure the localization capacity. The threshold is set to $0.2$ in our experiment. Average IoU is calculated over all the correctly predicted samples. Specifically, in few-shot image classification, $5$-way 11-shot is chosen for comparison, the average IoU is calculated over $10,000$ episodes. Table 1 and Table 2 show the quantitative comparison result. From the results, the proposed SFAM achieves the best performance in average IoU and localization accuracy, outperforming other methods by at least $2.2\%$.